OpenStudy (highlyeducated87):

Factoring: (-5x-2) - (-5x+2) First Answer gets medal it has to be right!!!

jimthompson5910 (jim_thompson5910):

Seems like something is missing here

jimthompson5910 (jim_thompson5910):

because if you simplify (-5x-2) - (-5x+2), you get -4

OpenStudy (highlyeducated87):

I kno

jimthompson5910 (jim_thompson5910):

I guess you could do -4 = -2*2 -4 = -1*4 but that seems too easy and I don't think that's what they meant